OpenStudy (anonymous):

Solve the equation. 6(2x - 3) + 4 = 16 - (-18)

OpenStudy (callisto):

6(2x - 3) + 4 = 16 - (-18) 6(2x) - 6(3) + 4 = 16 +18 12x - 18 + 4 = 34 12x - 14 = 34 12x = 48 x = ...?

OpenStudy (anonymous):

2

OpenStudy (callisto):

Nope. What is 48/12?

OpenStudy (anonymous):

oops wrng sorry i tyed wrong sorry!! its 4

OpenStudy (callisto):

Yes :)

OpenStudy (anonymous):

thnk you so much! :D

OpenStudy (callisto):

welcome!~